# Bridging and Modeling Correlations in Pairwise Data for Direct Preference Optimization

### v1

@@ 4 Experimental Setup
We conduct three downstream scenarios for a comprehensive evaluation, including question answering (QA), mathematical reasoning, and instruction following (IF). The detailed data statistics as well as the evaluation metrics are listed in Table 5 of Appendix A.1.
@@
For the QA and mathematical reasoning setup, we utilize Llama2-7B-base (Touvron et al., 2023) in our experiments. Successfully solving these tasks necessitates LLMs to possess domain-specific knowledge and engage in systematic, step-by-step reasoning to reach the ultimate answer. Therefore, following prior works (Chen et al., 2024a; b), we fine-tune Llama2-7B-base on the training set of ECQA (Aggarwal et al., 2021) and QASC (Khot et al., 2020) for QA, and fine-tune Llama2-7B-base on MetaMathQA (Yu et al., 2023) for mathematical reasoning. We denote the fine-tuned LLM as SFT and use it as the backbone for preference optimization. In line with prior research (Chen et al., 2024a; b), we construct preference pairs based on the training data, by using the ground truth as and the SFT model’s inference output as .
For the instruction following setup, we utilize Llama3-8B-base (AI@Meta, 2024) and Mistral-7B-Base (Jiang et al., 2023) in our experiments. Following the training pipeline of Zephyr (Tunstall et al., 2023) and SimPO (Meng et al., 2024), we train a base model on the UltraChat-200k dataset (Ding et al., 2023) to obtain an SFT model. Then, we use the SFT model as the starting point and perform preference optimization on the UltraFeedback dataset (Cui et al., 2023), where and are collected from LLMs of varying quality.
@@
In question answering, we adopt the test splits of ECQA (Aggarwal et al., 2021), QASC (Khot et al., 2020), OpenbookQA (Mihaylov et al., 2018), and StrategyQA (Geva et al., 2021) for evaluation. In mathematical reasoning, we conduct the evaluation on four challenge datasets including GSM8k (Cobbe et al., 2021), MATH (Hendrycks et al., 2021), MAWPS (Koncel-Kedziorski et al., 2016), and TabMWP (Lu et al., 2023). In instruction following, We assess our models using two of the most popular open-ended instruction-following benchmarks: AlpacaEval 2 (Li et al., 2023) and Arena-Hard v0.1 (Li et al., 2024). Both benchmarks evaluate the models’ versatile conversational abilities across a diverse set of queries. For each query, the evaluated model’s response and reference model’s response are compared head-to-head using an auto-evaluator. We use the officially recommended configurations111AlpacaEval: https://github.com/tatsu-lab/alpaca_eval. Arena-Hard v0.1: https://github.com/lm-sys/arena-hard-auto. during the evaluation.
@@
We compare our approach with various powerful offline preference optimization methods, including FIGA (Guo et al., 2023), DPO (Rafailov et al., 2024b), and DPO variants (IPO (Azar et al., 2024), KTO (Ethayarajh et al., 2024), ORPO (Hong et al., 2024), TDPO (Zeng et al., 2024), R-DPO (Park et al., 2024), and SimPO (Meng et al., 2024)). The training objectives of these methods are listed in Table 6.
@@ 5.1 Main results and ablations
As shown in Table 1, our model DPO-BMC achieves the best overall performance across all benchmarks on QA and math tasks. Notably, DPO-BMC outperforms DPO by 3.8 absolute points on QA tasks and by 1.3 points on math tasks. On instruction-following tasks, as shown in Table 2, DPO-BMC secures the highest length-controlled win rate, surpassing DPO by over 5 points across various settings. The length-controlled win rate (Dubois et al., 2024) is a debiased version of the win rate, significantly reducing length gameability in LLM-based auto-annotation. It can be seen that while achieving a higher raw win rate, DPO-BMC’s average response length is notably shorter than that of other baselines (a case study is presented in Table 8). This attribute is credited to the correlated preference data we constructed, which directs optimization towards critical desired behaviors rather than verbosity. Although FIGA aims to steer the model in a similar direction, it fails to improve performance over the SFT model, as it focuses solely on different tokens and employs fixed token-level rewards. These consistent and significant improvements underscore the robustness and effectiveness of our method.
@@
| Method | Llama3-Base (8B) | Mistral-Base (7B) | ||||||||
@@
Both key designs in BMC are crucial.
DPO-MC: Training using our proposed objective in Eq. 5 on the original preference data.
Our examination reveals several key findings: (1) enhancing the correlation between preference pairs remarkably boosts model performance; (2) even when using the same training preference data, our designed optimization objective outperforms DPO, highlighting its superior ability to model fine-grained correlations; and (3) combining our constructed data with our designed objective yields the best results, affirming the inseparability of the Bridging Phase and the Modeling Phase—bridging correlation and then modeling it.
Influence of .
Here we investigate the impact of the threshold in the DPO-BMC objective on model performance, as illustrated in Figure 3. Setting reduces our method to one that assigns fixed token-level rewards. The optimal value of varies across different tasks, highlighting the need for task-specific tuning. Additionally, a large value can degrade model performance due to excessively aggressive gradient updates on certain tokens, which negatively affects both QA and instruction-following tasks.
5.2 Quantitative analysis of bridging and modeling phase
To rigorously assess the effectiveness of the two pivotal phases in our framework, we segment the 60k training data of UltraFeedback into six equal-sized splits, ordered by increasing edit distance between the winning and losing responses. For each split, we also construct its corresponding pair data through our Bridging Phase. We then train four models—–(a) DPO, (b) DPO-MC, (c) DPO-BC, and (d) DPO-BMC—on each split, and report the length-controlled win rate on AlpacaEval 2 and the average gradient norm during training. All models are trained with identical hyperparameters to ensure consistency and comparability. It can be observed from Figure 4 that the Bridging Phase successfully decreases the edit distance between pairwise data through targeted modification, shifting the optimization process to concentrate on the most critical differences in preference data. The results underscore the efficacy of the Bridging Phase in consistently enhancing the performance across all splits, signifying its role in refining the training data for more focused learning. A notable observation is the behavior of the average gradient norm during the training of DPO, which exhibits an increasing trend as the edit distance between pairwise data enlarges. This trend highlights the sensitivity of DPO’s training process to individual data points, which can potentially lead to gradient imbalances and inefficient learning. However, the proposed learning objective in our Modeling Phase effectively mitigates these imbalances through a dynamic training process that adjusts according to the policy model’s confidence. This adaptive mechanism not only ensures that the model places greater emphasis on learning challenging distinctions but also reinforces the patterns that have already been acquired, culminating in a more nuanced and robust policy. Such an approach is crucial in achieving a balanced optimization landscape with diverse training data, leading to superior generalization and performance in downstream tasks.
In Figure 5, we present the KL divergence between the policy model trained with DPO, DPO-MC, DPO-BC, and DPO-BMC with identical hyperparameters and the reference model, measured on the winning responses from a held-out set of UltraFeedback during training. The results also validate our above analyses: (1) the Bridging Phase fosters tailored learning toward critical differences in preference data, potentially resulting in a larger KL divergence during training; (2) our meticulously designed loss function in the Modeling Phase effectively moderates the optimization intensity across diverse training data, thereby achieving a more controlled and steady KL divergence.
5.3 Quantitative analysis of credit assignment
We quantitatively compare the token-level and sequence-level credits assigned by DPO and DPO-BMC, assessing how well their final learned rewards align with preference labels on a held-out set of UltraFeedback.
Analysis on token-level reward.
Figure 6 depicts the token-level reward assignment for DPO and DPO-BMC on a response pair consisting of a preferred response and a dispreferred response . The reward of each token is computed as . From the figure, we observe that: (1) DPO assigns nearly uniform rewards to each token within a response, failing to differentiate the importance of tokens to the overall response quality; and (2) while DPO can identify and assign lower rewards to several erroneous tokens in the dispreferred response, such as “13” and assigning them small rewards, it struggles to discriminate subtle differences between the preferred and dispreferred responses. In contrast, DPO-BMC not only assigns relatively higher rewards to critical tokens like “descending order” but also precisely identifies the incorrect tokens inside the dispreferred response. Notably, within the incorrect span of “22, 13, 99”, DPO-BMC assigns a significantly low reward to the beginning token “22” while imposing lesser penalties on the subsequent ones. This pattern of reward assignment is more reasonable as it reflects the autoregressive nature of the generation, where subsequent mistakes are conditioned on previous errors. This example highlights the limitation of DPO in providing precise token-level preferences on sentence quality, and our method can effectively alleviate this issue. As illustrated, a generated sequence may contain a mix of correct and incorrect information, or even accurate and helpful responses marred by minor errors. To this end, our method provides a more effective and interpretable approach to discerning the subtle quality of these outputs.
Analysis on sequence-level reward.
For a rigorous comparison, we calculate the sequence-level DPO reward expression by (with the partition function excluded) and present the results in Table 4. The reward margin is determined by . Reward accuracy is defined as the percentage of preference pairs where the winning response achieves a higher reward than the losing response, i.e., . Our findings demonstrate that DPO-BMC outperforms DPO in terms of average reward margin and reward accuracy. This enhancement validates the superior ability of our method to discern subtle differences between preference pairs, thereby facilitating more effective generalization.
| Method | Avg. Reward Margin | Reward Accuracy (%) |
| DPO | 0.54 | 72.19 |
| DPO-BMC | 0.74 | 73.60 |
@@ 5.4 Versatility of our framework
Our BMC framework is versatile and can be applied to various DPO variants. Table 3 shows the performance comparison between the original XPO and XPO-BMC methods. Given that KTO only requires a binary signal to indicate whether an output is desirable or undesirable for a given input, and TDPO is based on token-level MDP, we exclude these from our consideration. The results indicate that XPO-BMC variants generally surpass their corresponding XPO methods across various tasks. Specifically, IPO-BMC, ORPO-BMC, R-DPO-BMC, and DPO-BMC achieve higher average accuracies in QA and Math, as well as superior LC scores on AlpacaEval 2 for IF. However, SimPO-BMC does not outperform SimPO on the IF task, showing a slight decrease from 21.3 to 19.7. These findings underscore the robustness and efficacy of our BMC approach in enhancing the performance of different DPO methods, affirming its potential for broad application.
@@ 6 Conclusion
Appendix A Appendix
@@ A.1 Data statistics and evaluation metrics used for experiments
| Task | Train / Test | Dataset | Number | Metric |
@@
| Test | AlpacaEval 2 | 805 | Win rate | |
@@ A.2 Prompt template for targeted modification
A.3 Implement details
To ensure a fair comparison, we conduct thorough hyperparameter tuning for all methods compared in our experiments.
@@ A.3 Implement details
A.4 Case Study

### v3

@@ 4 Experimental Setup
We conduct a comprehensive evaluation across three downstream scenarios, including question answering (QA), mathematical reasoning, and instruction following (IF). The detailed data statistics as well as the evaluation metrics are listed in Table 6 of Appendix A.1.
@@
For the QA and mathematical reasoning setup, we utilize Llama2-7B-base (Touvron et al., 2023) in our experiments. Dealing with these tasks necessitates LLMs to possess domain-specific knowledge and engage in systematic, step-by-step reasoning to reach the ultimate answer. Therefore, following prior works (Chen et al., 2024a; b), we fine-tune Llama2-7B-base on the training set of ECQA (Aggarwal et al., 2021) and QASC (Khot et al., 2020) for QA, and fine-tune Llama2-7B-base on MetaMathQA (Yu et al., 2024) for mathematical reasoning. We denote the fine-tuned LLM as SFT and use it as the backbone for preference optimization. In line with prior research (Chen et al., 2024a; b), we construct preference pairs based on the training data, by using the ground truth as and the SFT model’s inference output as . For the instruction following setup, we utilize Llama3-8B-base (Dubey et al., 2024) and Mistral-7B-Base (Jiang et al., 2023) in our experiments. Following the training pipeline of Zephyr (Tunstall et al., 2023) and SimPO (Meng et al., 2024), we train a base model on the UltraChat-200k dataset (Ding et al., 2023) to obtain an SFT model. Then, we use the SFT model as the starting point and perform preference optimization on the UltraFeedback dataset (Cui et al., 2023), where and are collected from LLMs of varying quality.
During our Bridging Phase, we utilize gpt-4-0125-preview for targeted modification to obtain , based on the prompt template in Appendix A.2. We also demonstrate in Table 4 that a less powerful open-source LLM, such as Llama3-70B-Instruct, can acquire comparable results. During our Modeling Phase, we list the implementation details in Appendix A.3 for reproducibility. A comprehensive cost analysis in Appendix B confirms that the computational overhead introduced by our BMC pipeline is minimal.
@@
In question answering, we adopt the test splits of ECQA (Aggarwal et al., 2021), QASC (Khot et al., 2020), OpenbookQA (Mihaylov et al., 2018), and StrategyQA (Geva et al., 2021) for evaluation. In mathematical reasoning, we conduct the evaluation on four challenge datasets including GSM8k (Cobbe et al., 2021), MATH (Hendrycks et al., 2021), MAWPS (Koncel-Kedziorski et al., 2016), and TabMWP (Lu et al., 2023). In instruction following, We assess our models using two of the most popular open-ended instruction-following benchmarks: AlpacaEval 2 (Li et al., 2023) and Arena-Hard v0.1 (Li et al., 2024). Both benchmarks evaluate the models’ versatile conversational abilities across a diverse set of queries. For each query, the evaluated model’s response and the reference model’s response are compared head-to-head using an auto-evaluator. We use the officially recommended configurations111AlpacaEval: https://github.com/tatsu-lab/alpaca_eval. Arena-Hard v0.1: https://github.com/lm-sys/arena-hard-auto. during the evaluation.
@@
We compare our approach with various powerful offline preference optimization methods, including FIGA (Guo et al., 2024), DPO (Rafailov et al., 2024b), and DPO variants (IPO (Azar et al., 2024), ORPO (Hong et al., 2024), R-DPO (Park et al., 2024), and SimPO (Meng et al., 2024)). The training objectives of these methods are listed in Table 7. Besides, we include two additional baselines: (1) DPO (CW): enhancing pairwise data correlation by prompting the SFT model to Continue Writing a prefix of the winning response to generate the losing one; (2) DPO (EW): leveraging an off-the-shelf LLM for External Weighting of token-level reward (Lee et al., 2024), where LLM scores each token in the winning and losing responses based on how much it improves or decreases the overall quality.
@@ 5.1 Main results and ablations
As presented in Table 1, our model DPO-BMC consistently achieves state-of-the-art results across all evaluated QA and math benchmarks. Specifically, DPO-BMC outperforms DPO by 3.8 absolute points on QA tasks and by 1.3 points on math tasks. On instruction-following tasks (Table 2), DPO-BMC secures the highest length-controlled win rate, surpassing DPO by over 5 points across various settings, with even greater gains for larger base models (Appendix E). The length-controlled win rate (Dubois et al., 2024) serves as a robust metric that mitigates the effects of length bias, thereby providing a more reliable evaluation of LLM-based auto-annotation. Notably, DPO-BMC generates responses that are significantly more concise than other baselines. As highlighted in Table 2, the average response length of DPO-BMC and DPO-BC is approximately 75% of that produced by DPO and DPO-MC. This attribute of length normalization is credited to the correlated preference data we constructed, which directs optimization towards critical desired behaviors rather than verbosity. A case study in Table 12 further underscores the effectiveness and robustness of our approach.
@@
| Method | Llama3-8B-Base | Mistral-7B-Base | ||||||||
@@
Both key designs in BMC are crucial.
DPO-MC: Training using our proposed objective in Eq. (4) on the original preference data.
Our examination reveals several key findings: (1) DPO (CW), the “Continue Writing” approach, slightly underperforms standard DPO, as it introduces superficial correlations that fail to capture the nuanced, task-specific alignments essential for effective optimization. In contrast, our Bridging Phase explicitly enhances informative correlations—elucidate fine-grained distinctions between desired and undesired behaviors through token-level variations. This targeted focus significantly improves model performance; (2) Even when leveraging identical training preference data, our designed optimization objective consistently outperforms both DPO and DPO (EW), highlighting its superior ability to model fine-grained correlations based on the dynamic of the policy model’s confidence; and (3) Combining our constructed data with our designed objective yields the best results, affirming the inseparability of the Bridging Phase and the Modeling Phase.
| Data Synthesis | Training Data | QA | Math | IF |
| (ours) | 65.1 | 49.6 | 22.4 | |
| 64.3 | 49.2 | 19.8 | ||
| 64.6 | 48.7 | 18.9 | ||
| 63.9 | 48.6 | 17.6 |
Influence of data synthesis method.
Table 3 shows the effects of various data synthesis strategies during the Bridging Phase. When generating without referring to , LLMs potentially make erroneous modifications that misalign with the intended target, leading to a performance drop. An alternative approach that attempts to enhance the correlation between winning and losing responses is to degenerate to and utilize as the preference pair. However, this approach also falls short, likely because LLMs are primarily trained to generate high-quality data, making it challenging for them to generate low-quality outputs that mimic the nuanced errors of losing responses. Semantic similarity analysis using the all-mpnet-base-v2 embedding model222https://huggingface.co/sentence-transformers/all-mpnet-base-v2 supports this, showing a high score of 0.88 for but only 0.73 for .
Influence of data modification proportion.
Figure 3 illustrates the impact of data modification proportions during the Bridging Phase on performance. Increasing modifications from 0% to 20% yields the most substantial gains, highlighting the effectiveness of enhancing pairwise preference correlations. Performance plateaus beyond 80% modifications, indicating that extensive changes are beneficial but not essential, offering flexibility under computational or data constraints. These results demonstrate the scalability and adaptability of our framework for diverse applications.
Influence of LLMs for targeted modification.
Table 4 explores the influence of diverse LLMs for targeted modification. Notably, substituting the gpt-4-0125-preview model with a less powerful yet open-source alternative, such as Llama3-70B-Instruct, yields comparable performance while significantly surpassing vanilla DPO. This finding underscores the adaptability of our method to varying levels of model sophistication, thereby reducing dependence on commercial LLMs without significant impact on final model performance.
Influence of .
We conduct an ablation study to examine the influence of the threshold in the DPO-BMC objective on model performance, as shown in Figure 4. Setting reduces our method to one that assigns fixed token-level rewards, leading to suboptimal accuracy. As increases, the model performance improves, with the optimal setting observed around . However, further increasing results may degrade model performance due to excessively aggressive gradient updates on certain tokens. Notably, across all tested values of , our method consistently outperforms the DPO baseline, indicating its robustness and effectiveness in stabilizing the learning process.
| Method | LLM for Targeted Modification | QA | Math | IF |
| SFT | – | 56.9 | 47.6 | 7.5 |
| DPO | – | 61.3 | 48.3 | 16.0 |
| DPO-BMC | Llama3-70B-Instruct | 64.6 | 49.4 | 21.8 |
| DPO-BMC | gpt-4-0125-preview | 65.1 | 49.6 | 22.4 |
5.2 Quantitative analysis of bridging and modeling phase
To rigorously assess the effectiveness of the two pivotal phases in our framework, we segment the 60k training data of UltraFeedback into six equal-sized splits, ordered by increasing edit distance between winning and losing responses. For each split, we also construct its corresponding pair data through our Bridging Phase. We then train four models—–(a) DPO, (b) DPO-MC, (c) DPO-BC, and (d) DPO-BMC—on each split based on Llama3-8B, with identical hyperparameters to ensure comparability. As shown in Figure 5, the Bridging Phase successfully decreases the edit distance between pairwise data through targeted modification, shifting the optimization process to concentrate on the most critical differences in preference data. This phase consistently enhances performance across all splits by refining training data for more focused learning. Another notable observation is the average gradient norm during DPO training increases as the edit distance between pairwise data enlarges, reflecting the sensitivity of DPO’s training process to individual data points and potential gradient variance. Our proposed Modeling Phase mitigates the variance by dynamically adjusting the training process based on the policy model’s confidence. This adaptive mechanism prioritizes challenging distinctions while reinforcing learned patterns, promoting a balanced optimization landscape with diverse training data (See Appendix D for further analysis).
5.3 Quantitative analysis of credit assignment
We compare the token-level and sequence-level credits assigned by DPO and DPO-BMC, assessing how well their final learned rewards align with preference labels on a held-out set of UltraFeedback.
Analysis on token-level reward.
Figure 6 depicts the token-level reward assignment for DPO and DPO-BMC on a response pair consisting of a winning response and a losing response . The reward of each token is computed as . From the figure, we observe that: (1) DPO assigns nearly uniform rewards across tokens, failing to differentiate the importance of tokens to the overall response quality; and (2) although DPO can identify and assign lower rewards to several erroneous tokens in the losing response (e.g., “13”), it struggles to capture subtle distinctions between the winning and losing responses. In contrast, DPO-BMC assigns higher rewards to critical tokens (e.g., “descending order”) and effectively penalizes incorrect tokens in the losing response. These results demonstrate DPO’s limitations in providing precise token-level preferences on sentence quality, and our method can effectively alleviate this issue.
Analysis on sequence-level reward.
For a rigorous comparison, we calculate the sequence-level DPO reward expression by . The reward margin is determined by . Reward accuracy is defined as the percentage of preference pairs where the winning response achieves a higher reward than the losing response, i.e., . Our findings show that DPO-BMC outperforms DPO in terms of average reward margin (0.74 vs. 0.54) and reward accuracy (73.60 vs. 72.19). This enhancement validates our method’s superior ability to discern subtle differences between preference pairs, enabling more effective generalization.
@@ 5.4 Versatility of our framework
Our BMC framework demonstrates versatility and can be seamlessly integrated with various DPO variants. As shown in Table 5, the XPO-BMC methods consistently outperform their corresponding XPO baselines across a diverse set of tasks, including QA, Math, and Instruction Following (IF). For instance, IPO-BMC achieves a significant improvement in QA accuracy (64.1 vs. 60.6) and IF score (15.7 vs. 13.4) compared to IPO. Similarly, ORPO-BMC, R-DPO-BMC, SimPO-BMC, and DPO-BMC exhibit higher performance in QA and Math, alongside notable gains in IF, such as R-DPO-BMC improving the IF score from 17.1 to 20.0 over R-DPO. These results highlight the robustness of our framework in enhancing task-specific performance across various settings, reaffirming its potential as a generalizable enhancement to existing DPO methodologies.
@@ 6 Conclusion
Appendices
Appendix A Detailed Experimental Setup
@@ A.1 Data statistics and evaluation metrics used for experiments
| Task | Train / Test | Dataset | Number | Evaluation Metric |
@@
| Test | AlpacaEval 2 | 805 | Win rate against GPT-4 Turbo | |
@@ A.2 Prompt template for targeted modification
A.3 Implementation details
Our implementation is based on the alignment-handbook repo333https://github.com/huggingface/alignment-handbook using 4×A800 GPUs. To ensure a fair comparison, we conduct thorough hyperparameter tuning for all methods compared in our experiments.
@@ A.3 Implementation details
Appendix B Cost Analysis
B.1 Cost of Bridging Phase
The Bridging Phase, responsible for synthesizing pseudo-winning responses, operates exclusively offline, meaning it incurs no runtime cost during model training. The data synthesis process is designed to be efficient, as it does not require iterative computations or model updates.
For context, we estimated the budget for data synthesis using the gpt-4-0125-preview API, based on the API’s pricing of $0.01 per 1K input tokens and $0.03 per 1K output tokens. Table 9 lists the breakdown of the estimated costs for our three evaluated tasks, which demonstrates that this is a manageable expenditure.
| Task | # of Samples | Avg. Input Token Length | Avg. Output Token Length | Cost ($) |
| QA | 15,732 | 206 | 25 | 44.21 |
| Math | 40,000 | 429 | 47 | 228.00 |
| IF | 61,135 | 728 | 235 | 876.06 |
Can an Open-Source LLM be Utilized as an Alternative?
In Table 4, we explore the impact of LLMs on targeted modifications during the Bridging Phase. Our findings indicate that substituting the gpt-4-0125-preview model with a less powerful yet open-source alternative, such as Llama3-70B-Instruct, yields comparable performance while significantly surpassing vanilla DPO. The Llama3-70B-Instruct model can be deployed on only 2 NVIDIA-3090 GPUs, with the option to further reduce hardware requirements through low-bit quantization444https://github.com/ollama/ollama. This provides an economical alternative for our Bridging Phase without compromising performance. Numerous studies have highlighted the superior text modification capabilities of LLMs. For example, LLMs have been effectively employed in synthesizing high-quality data (Wang et al., 2023; Liu et al., 2024b). Additionally, Ji et al. (2024) show that LLMs can transform initial outputs from upstream models into more helpful and benign responses, thereby aligning generated content with human intentions. In conclusion, our framework demonstrates robustness in leveraging diverse LLMs for targeted modifications, confirming its adaptability and effectiveness.
B.2 Cost of Modeling Phase
Our Modeling Phase adds minimal computational overhead compared to vanilla DPO. Specifically:
Token Difference Identification: Using a dynamic programming algorithm (edit distance) to identify differing tokens between the pseudo-winning and losing responses. This is a lightweight operation and introduces negligible runtime cost.
Reward Weighting Calculation: We calculate a weighting factor based on the policy model’s probability of the identified tokens, which is already computed in the standard DPO setup. Because we halt gradient backpropagation for the weighting factor, this operation does not introduce additional computational costs.
Table 10 demonstrates the comparison of the training times between DPO and DPO-BMC on 4×A800 GPUs, illustrating that DPO-BMC increases training time by less than 1% across all evaluated tasks.
| Task | Base Model | Runtime of DPO (s) | Runtime of DPO-BMC (s) | Increase (%) |
| QA | Llama2-7B | 2,831 | 2,850 | 0.67% |
| Math | Llama2-7B | 9,586 | 9,641 | 0.57% |
| IF | Llama3-8B | 16,179 | 16,318 | 0.86% |
Overall, these results validate that the computational overhead introduced by BMC is minimal, and the approach is highly efficient in terms of runtime, making it practical for real-world applications without significantly increasing resource requirements.
Appendix C Gradient analysis
For a mechanistic understanding of our method, we examine the gradients of the loss function in Eq. (2) and in Eq. (4). Their gradients with respect to the parameters can be written as:
where .
where .
In contrast to vanilla DPO, which emphasizes sequence-level optimization exclusively, our proposed method integrates both sequence-level and token-level perspectives. (1) At the sequence level, we promote preferred completions while penalizing those that are disfavored. (2) At the token level, we further refine the rewards of critical desired and undesired tokens of and , respectively. This dual consideration ensures that both the overall sequence structure and the critical token choices are optimized for the desired outcome.
Appendix D KL Divergence Analysis During Training
In Figure 9, we present the KL divergence between the policy model trained with DPO, DPO-MC, DPO-BC, and DPO-BMC with identical hyperparameters and the reference model, measured on the winning responses from a held-out set of UltraFeedback during training. The results also validate our analyses in §5.2: (1) the Bridging Phase fosters tailored learning toward critical differences in preference data, resulting in more efficient and “sharp” training with a larger KL divergence; (2) our meticulously designed loss function in the Modeling Phase effectively moderates the optimization intensity across diverse training data, thereby achieving a more controlled and steady KL divergence.
Appendix E Experiments on larger base models
We conducted additional experiments using the more capable Qwen2.5-14B-Base model (Team, 2024). As shown in Table 11, our proposed DPO-BMC method delivers even greater performance improvements with this model, achieving a remarkable gain of +8.8 on AlpacaEval 2 and +7.3 on Arena-Hard compared to standard DPO. These results highlight the effectiveness of DPO-BMC scales with model capability, underscoring its potential to deliver even larger gains when applied to more powerful baseline models.
| Method | Base Model | LC (%) of AlpacaEval 2 | WR (%) of Arena-Hard |
| SFT | Mistral-7B-Base | 8.1 | 2.2 |
| DPO | Mistral-7B-Base | 15.1 | 13.6 |
| DPO-BMC | Mistral-7B-Base | 20.8 (+5.7) | 17.6 (+4.0) |
| SFT | Llama3-8B-Base | 7.5 | 2.6 |
| DPO | Llama3-8B-Base | 16.0 | 17.6 |
| DPO-BMC | Llama3-8B-Base | 22.4 (+6.4) | 18.1 (+0.5) |
| SFT | Qwen2.5-14B-Base | 14.1 | 11.2 |
| DPO | Qwen2.5-14B-Base | 36.3 | 33.6 |
| DPO-BMC | Qwen2.5-14B-Base | 45.1 (+8.8) | 40.9 (+7.3) |
Appendix F Case Study